(To monitor the crewmember's sleep/wake patterns and light exposure, Garrett wears a special Actiwatch device which measures the light levels encountered by him as well as his patterns of sleep and activity throughout the Expedition. The log entries are done within 15 minutes of final awakening for seven consecutive days.)
Also before breakfast, having reached the FD15 (Flight Day 15) mark in his flight, Reisman undertook his first session with the NASA/JSC experiment NUTRITION w/Repository, collecting blood and urine samples. (Acting as operator and CMO (Crew Medical Officer), CDR Whitson performed phlebotomy on Garrett, i.e., drew blood samples (from an arm vein) which were first allowed to coagulate in the Repository, then spun in the HRF RC (Human Research Facility/Refrigerated Centrifuge) and finally placed in MELFI (Minus-Eighty Laboratory Freezer for ISS). The RC was later powered off after a temperature reset to limit wear on the compressor, and cleaned. Garrett's urine samples were also placed in the MELFI. Background: NUTRITION is the most comprehensive in-flight study done by NASA to date of human physiologic changes during long-duration space flight; this includes measures of bone metabolism, oxidative damage, nutritional assessments, and hormonal changes. The Clinical Nutritional Assessment profile currently required on all U.S. Astronauts collects blood and urine samples preflight and postflight. NUTRITION expands this protocol by also capturing in-flight samples and an additional postflight sample. Furthermore, additional measurements are included for samples from all sessions, including additional markers of bone metabolism, vitamin status, and hormone and oxidative stressor tests. The results will be used to better understand the impact of countermeasures (exercise and pharmaceuticals) on nutritional status and nutrient requirements. The Clinical Nutritional Assessment profile (MR016L), first started on two Mir crewmembers and then on all ISS US crews, nominally consists of two pre-flight and one post-flight analysis of nutritional status, as well as an in-flight assessment of dietary intake using the FFQ (Food Frequency Questionnaire). The current NUTRITION project expands MR016L testing in three ways: Addition of in-flight blood & urine collection (made possible by MELFI), normative markers of nutritional assessment, and a return session plus 30-day (R+30) session to allow evaluation of post-flight nutrition and implications for rehabilitation.)
In preparation for today's ATV1 'Jules Verne' Demo Day 1, Peggy Whitson & Yuri Malenchenko -
closed the Lab science window shutters as protection against Russian thruster contamination,
set up the KL-154 'Klest' television equipment in the SM (Service Module) with the ATV TV control console (BRTK-PU),
configured the designated A31p laptop in the FGB for transmitting analog & digital video of the activities from the RS (Russian Segment) from the MPEG-2 (Moving Pictures Expert Group 2) encoder, downlinked via U.S. OpsLAN and Ku-band in 'streaming video' packets, and
deactivated the A31p after the rendezvous demo activities.
ISS attitude control authority was handed over to RS MCS (Motion Control System) thrusters at 9:30am EDT for maneuvering the station to ATV1 rendezvous attitude at 10:04am. (Rendezvous attitude: LVLH (Local Vertical Local Horizontal) +XVV (+x-axis in velocity vector), YPR (Yaw, Pitch, Roll) = 0,0,0. Later, after returning to TEA (Torque Equilibrium Attitude) at ~1:37pm, control was handed back to U.S. momentum management at 2:12pm.)
After preparing for the ATV1 rendezvous demo in the SM, Malenchenko & Whitson activated and checked out the MBRL/PCE (Proximity Communications Equipment) Space-to-Space Radio hardware of the ASN-M satellite navigation system, then monitored approach, station-keeping and subsequent escape maneuvers of the remote-controlled ATV1, recording times and performance data. (The Demo Day 1 program involved Jules Verne closing in on ISS from ~39km behind (at 10:18am EDT) to 15.5. km (by 11:02am) and arriving at the S2 holding point at a distance of ~3500m behind ISS at ~11:48am for subsequent testing of HOLD, RETREAT, ESCAPE and other critical commands through PCE without actually executing them. Afterwards, the actual Escape maneuver of 4 m/sec delta-V was to take the vehicle away from and below the station at ~1:29pm.)
Afterwards, Peggy & Yuri were scheduled (~2:30pm) for a debriefing teleconference with the ground to review the ground/crew communications and determine if anything needs to be changed before Demo Day 2.
FE-2 Reisman meanwhile performed the periodic offloading of the Lab CCAA (Common Cabin Air Assembly) dehumidifier's condensate tank, filling CWC (Contingency Water Container) #1062 with the collected water slated for processing. No samples required this time. (Estimated offload time before termination (leaving ~6 kg in the tank): ~35 min.)
In the COL (Columbus Orbital Laboratory), Garrett also checked out the EMCS (European Modular Cultivation System) applications on the laptop computer (ELC) in the newly relocated US ER3 (EXPRESS Rack 3) which he connected yesterday for ground-commanded power-up. (The EMCS rack contains a rotating centrifuge that can support a wide range of small plant and animal experiments under partial gravity conditions.)
Yuri Malenchenko had another hour set aside for his end-of-increment cleanup and departure preparations. (Instructions on packing of return items and a keep vs. trash list were uplinked for assisting Yuri and Peggy in their preparing for their return in the severely downmass-limited Soyuz Descent Module. Trashed items will be stowed in the Orbital Module, to be separated along with the Instrumentation/Propulsion Module prior to atmospheric entry.)
Afterwards, the FE-1 completed the routine maintenance of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM, including ASU toilet facilities systems/replaceables. (Regular daily SOZh maintenance consists of replacement of the KTO & KBO solid waste containers, replacement of an EDV-SV waste water and EDV-U urine container, replacement of the KOV EDV for the Elektron-intended water, and processing U.S. condensate water as it becomes available in a filled CWC from the Lab humidifier.)
Whitson conducted the periodic (monthly) CSA-CP (Compound Specific Analyzer-Combustion Products) maintenance/checkout, today on all four units. (The CSA-CP is a passive cabin atmosphere monitor that provides quick response capability during a combustion event (fire). Its collected data are stored on a logger. Peggy changed out the battery on the prime unit, then zero-calibrated all instruments (to eliminate drift in the combustion sensors). Following zero calibration, the backup units were stowed in the Node, along with the sampling pump, while the prime unit was deployed at the SM Central Post.)
Yuri performed another run with the Russian "Uragan" (hurricane) earth-imaging program, using the Nikon D2X digital camera through SM windows 6-9. (Targets uplinked for today were the country of Columbia (natural environment) and the city of Bogotá.)
Working off his 'available time' discretionary task list, Malenchenko also performed the regular daily checkup on the Japanese experiment GCF-JAXA (Granada Crystallization Facility) in the Russian TBU incubator,
The crewmembers performed their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2), TVIS treadmill (CDR, FE-1/full time, FE-2), and RED resistive exercise device (CDR, FE-2).
Afterwards, the CDR transferred the crew's exercise data file to the MEC (Medical Equipment Computer) laptop for downlinking, as well as the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).
At ~8:05am, the crewmembers held their regular WPC (Weekly Planning Conference) with the ground, discussing next week's "Look-Ahead Plan" (prepared jointly by MCC-H and TsUP/Moscow timeline planners) via S-band/audio, reviewing the monthly calendar, upcoming activities, and any concerns about future on-orbit events.
ATV Update (Flight Day 20): Yesterday (3/28), the ATV1 performed the following three maneuvers, all of which were reported to be nominal:
TB1: start 6:48:53pm EDT, delta-V: 5.976 m/s
TB2: start 7:38:13pm, delta-V: 6.947 m/s
TB3: start 8:30:32pm, delta-V: 3.054 m/s.
ATV-CC/Toulouse completed an RDS (Russian Docking System) checkout. The preliminary results were nominal; however a detailed evaluation is being performed by the ATV-CC Engineering Support Team. In addition, the ISS crew completed the third ATV rendezvous on-board training session yesterday. Due to a problem with the simulator software on the crew's laptop, a subset of the session will be repeated on 3/30.
ATV-CC continues to see higher than expected power consumption from the pressurized module shell heaters and preliminary investigations indicate the problem is due to loose or missing thermal blankets. MART (Multi-lateral Anomaly Resolution Team) meetings were held on 3/27 & 3/28 and concluded that there will be no impacts to docking or crew ingress, and no impact to undocking has been identified.
Weekly Science Update (Expedition Sixteen -- Week 23)
ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Radiation measurements continue to be performed in the PIRS module. Radiation measurements continue to be performed in the PIRS module.
BCAT-3 (Binary Colloidal Alloy Test 3): Reserve.
CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): Reserve.
CFE (Capillary Flow Experiment): Reserve.
CSI-2/CGBA (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus): Two sides of first Silicate Garden complete with students analyzing data in their classrooms currently, and two Petri Dish runs complete but with no results. The Petri Dish installed during 1JA joint ops contains two small tomato plants as seen in the downlinked images with a possible third new one. This Petri Dish will remain in place until 1J when it will be returned for analysis of plant material.
CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.
CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): In progress.
EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.
ELITE-S2 (Elaboratore Immagini Televisive - Space 2): Planned.
EPO (Educational Payload Operations): Reserve.
ETD (Eye Tracking Device): Last session (4th) is currently planned just before Soyuz docked Ops.
EuTEF (European Technology Exposure Facility): The EuTEF platform power feeders have been successfully de-activated/re-activated in conjunction with the EVA-3 and EVA-5 on 3/18 and 3/23, respectively. On 3/24, a graceful shutdown was performed after off-nominal and incoherent telemetry readings. EuTEF is nominal since then. - DEBIE-2: the instrument showed regular link errors. Currently powered off; - DOSTEL: On-going science acquisition; - EuTEMP: Currently inactive as planned; - EVC: after a successful commissioning, we encountered some difficulty to get nominal images acquisition. Under investigation if the problem comes from EVC itself or from the Columbus High Rate Multiplexer; Currently powered off; - EXPOSE: On-going science acquisition; - FIPEX: FIPEX showed Link Errors and was further troubleshooted this week; tests are on-going and basic science measurements are collected; - MEDET: On-going science acquisition; - PLEGPAY: The first long-duration (100 hrs) experiment has been started on 3/26; Currently powered on; - TRIBOLAB: in Stand-by mode, awaiting to start science acquisition after ATV docking on 4/3; Currently powered on.
Fluid Science Laboratory (FSL): The FSL Facility is awaiting further troubleshooting after 1J/A departure.
GEOFLOW: On 3/25, the FSL VMU (Video Management Unit) Optical Fibre has been tested and is fully functional. The FSL commissioning will continue in the next coming weeks. FSL LAN and MIL bus repair activities have been postponed until after 1J/A mission. Start of GEOFLOW is pending further FSL troubleshooting/commissioning activities.
IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Blood and urine samples of first session have been downloaded on STS-123 (1J/A). Second session is currently planned during Soyuz docked Ops.
InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): In progress.
Integrated Immune: 'Peggy, thank you for unstowing Garrett's saliva kit and having it ready for him to begin his collections. Garrett, we appreciate your efforts during your early increment Integrated Immune session.'
KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.
LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Complete.
MISSE (Materials ISS Experiment): Ongoing.
MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 'Pirs'.
MSG-SAME (Microgravity Science Glovebox): Complete.
NOA-2 (Nitric Oxide Analyzer): Planned.
PMDIS (Perceptual Motor Deficits in Space): Complete.
SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.
SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): 'Peggy, we have placed on your task list next week the next Sleep download activity. During this activity you will be changing the battery in your Actiwatch as well as downloading and initializing 1J/A FE-2 and your Actiwatches. You will also change the battery and initialize an Actiwatch for the South-Korean SFP (KARI Space Flight Participant). Thanks for your participation.'
SOLAR (Solar Monitoring Observatory): Anomaly on platform Sun pointing mode under further investigation: no science acquisition possible so far. However, calibration runs of SOLSPEC spectral lamps have been performed on 3/26 and 3/27. On 3/27, SOLAR went unexpectedly in stand-by mode, but has been fully recovered. The calibration run on 3/27 will have to be repeated.
SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): In progress.
Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.
TRAC (Test of Reaction & Adaptation Capabilities): Planned.
ULTRASOUND: 'Peggy, thank you for moving and checking out the Ultrasound - we were VERY excited to see if functioning nominally!'
WAICO #1 (Waving and Coiling of Arabidopsis Roots at Different g-levels): The 4 ECs (Experiment Containers) from Rotor A (0-g condition) transfer from BIOLAB TCU#1 (Temperature Control Unit 1) to MERLIN (Microgravity Experiment Research Locker Incubator II) cooler unit has been successfully performed on 3/24. The MERLIN temperature was monitored on a daily basis and was very stable. These 4 ECs have been downloaded on STS-123 (1J/A) and currently shipped to Germany. The 4 ECs of Rotor B are yet to be retrieved by the crew from the blocked Rotor B. BIOLAB anomalies encountered during WAICO-Run#1 imply major impact on science return.
CEO (Crew Earth Observation): Ongoing.
CEO photo targets (for discretionary picture taking) uplinked for today were Spider Impact Crater, W Australia (this 13-km-wide crater is old (>570 million years). A comprehensive search of the CEO database shows no handheld images), Hyderabad, India (looking immediately left for this important city of 6.7 million. Visual cue is the dark line of the river that runs through the city. Trying to shoot the diffuse urban boundary), Somalia Coast (CEO observers are documenting longer term changes of vegetation in this semi desert. Shooting a swath following the coast. Vegetation seems to green and die off in concert with multi-year El Niños), and Kwanza Basin, Angola (a mapping swath under track was requested to document new economic development as oil wealth streams into Angola. Looking for linear features such as railroads, highways, and power lines).